sanju:

A 4.0 kg ball is swung at a constant speed in a vertical circle of radius 2.5 m. If the ball completes one revolution in 3.0 s, what is the tension in the rope at the top of the circle?

XxXNessalulbaddieXxX:

uhhhh @smokeybrown

SmokeyBrown:

This is another question involving centripetal force, so I assume you are still familiar with the equation from before, involving mass, velocity, and radius. This time, since the circle is vertical, you will also have to account for the affect of gravity on the ball. The question also specifies that you must find the tension in the rope "at the top of the circle", and I am not sure what affect that will have. It is a centripetal force question, with some extra parts. Unfortunately, I am not sure about how to deal with those extra parts. Perhaps we can try tagging someone with more specific physics knowledge?

Vocaloid:

draw the free body diagram of the situation. at the top of the circle, the tension and the gravity vectors point toward the center of the circle. the sum of these two vectors is your centripetal force Fc  gravity is simply Fg = mg Tension is Ft (which you'll solve for) and Fc = mv^2 / r from here, mg + Ft = mv^2/r m and r are given. velocity can be determined by circumference/time, or 2pi*r/t